Q/Let X {1,2,3, 4} and T = {0,X,{1}, {1,2). {1,2,3), {1,3}, {1,3, 4}, where
(X,T) is topological space. If A = {1,3,4}, B = {2,3} then find A, B°, ext (A), ext(B), b (A), b(B)
1
Expert's answer
2020-11-13T15:57:09-0500

"B^o" = union of subsets of B open in X. But they are only "\emptyset." "A^o=" "\{1,3,4\}" since A is open.

ext(B)= union of open sets disjoint from B = "\{1\}." ext(A)="\emptyset."

b(B)= "\overline{B} \setminus B^o=\overline{B}." Now "\{1\}^c=\{2,3,4\}". So the latter is closed and contains B. Also B is not closed since "B^{c}=\{1,4\}" is not open. Hence "\{2,3,4\}" is the smallest closed set containing B. Hence "b(B)=\overline{B}=\{2,3,4\}." "\overline{A}=X" since only other option is "A" itself and "A^c=\{2\}" not open, hence A not closed. Hence b(A)=X"\setminus A^o=X\setminus A=\{2\}".